# SiamSeg: Self-Training with Contrastive Learning for Unsupervised Domain Adaptation Semantic Segmentation in Remote Sensing

### v2

@@ IV-D2 Visualization Results
From Tables I, II, III, IV and V , we present the experimental results of SiamSeg across five cross-domain remote sensing image semantic segmentation tasks. In this section, we further visualize the model’s predictions to validate the outstanding performance of SiamSeg. Additionally, to explore the effectiveness of Contrastive Learning (C.L.), we visualize and compare the predictions of SiamSeg without C.L. (SiamSeg w.o. C.L.).
As shown in Fig. 4, we display the visualization results for the Potsdam and Vaihingen datasets, while Fig. 5 illustrates the results for the LoveDA Rural to Urban task. A comparison of these visualizations reveals that, overall, both models perform well. However, in certain local details, such as boundary recognition of buildings and roads and the smoothness of class differentiation, SiamSeg demonstrates a clear advantage over the model that does not incorporate contrastive learning. This indicates that the introduction of C.L. enhances the model’s understanding of image semantics and improves the extraction of network features.
In Fig. 4, we observe that in the task from Potsdam IRRG to Vaihingen IRRG, SiamSeg excels in recognizing the ”Building” and ”Tree” categories, while SiamSeg w.o. C.L. performs relatively poorly. This finding aligns with the experimental results presented in Table I. In the Potsdam RGB to Vaihingen IRRG task, the predictions on the left indicate that SiamSeg performs exceptionally well in recognizing ”Impervious Surface” and ”Low Vegetation,” particularly demonstrating high accuracy in predicting ”Impervious Surface” and its boundaries, consistent with the results in Table II.
Furthermore, in the task from Vaihingen IRRG to Potsdam IRRG, SiamSeg achieves outstanding performance in the ”Tree” and ”Low Vegetation” categories, accurately identifying small road structures, especially the roads within buildings in the left image. This observation corresponds with the experimental results in Table III. Although SiamSeg’s performance in the ”Clutter” category still shows some limitations, it exhibits significant improvement over the model without C.L. in this challenging category, further validating the effectiveness of the proposed contrastive learning approach. In the Vaihingen IRRG to Potsdam RGB task, SiamSeg excels in recognizing the ”Impervious Surface” category and demonstrates accurate identification in the ”Clutter” category, a finding that closely matches the experimental results in Table IV.
Observing Fig. 5, it is evident that SiamSeg generates precise predictions in cross-domain remote sensing image semantic segmentation tasks, particularly in the ”Building” and ”Barren” categories, where it delineates clear semantic boundaries. Notably, some regions classified as ”Road” in the predictions are labeled as ”Background” in the ground truth (GT). However, we do not consider this a misclassification by the model. Upon visual inspection, we find that these predicted ”Road” areas, although marked as ”Background” in the ground truth labels, exhibit characteristics typical of ”Road” from a semantic understanding perspective. This suggests that SiamSeg leverages more than just a data-driven learning mechanism; it demonstrates a stronger categorical understanding through deep feature induction and extraction, resulting in semantically plausible predictions in certain scenarios.
@@
Figure 4: Visualization of results on Potsdam and Vaihingen datasets. For each task, we provide 3 cases containing 4 images. From left to right, the 4 images are respectively target images, prediction of SiamSeg without contrastive learning, prediction of SiamSeg and ground truth. The cross-domain tasks from top to bottom are Potsdam IRRG to Vaihingen IRRG, Potsdam RGB to Vaihingen IRRG, Vaihingen IRRG to Potsdam IRRG and Vaihingen IRRG to Potsdam RGB. By looking at the visualization results, we can find that SiamSeg with Contrastive learning has better segmentation performance. The categories represented by the different colors are listed at the bottom of the picture with their names and colors.
Figure 5: Visualization of results on LoveDA datasets. We conduct one task which is Rural to Urban. We provide the visualization results on LoveDA dataset. Since images in the testing dataset do not have annotations, we display the results of images in the validation dataset. For this task we provide 1 case containing 4 images. From left to right, the 4 images are respectively target images, prediction of SiamSeg without contrastive learning, prediction of SiamSeg and ground truth. The categories represented by the different colors are listed at the bottom of the picture with their names and colors.
@@ IV-D3 Ablation Studies
Using the performance of SiamSeg and SiamSeg without Contrastive Learning (C.L.) in the task from Potsdam IRRG to Vaihingen IRRG, as shown in Table I, we can clearly observe that SiamSeg outperforms SiamSeg w.o. C.L. across all categories after the introduction of C.L. This indicates that C.L. effectively enhances the model’s ability to perceive features across different categories. The additional supervisory signals provided by C.L. enable the model to learn image features more profoundly, significantly improving performance in cross-domain semantic segmentation tasks. The introduction of C.L. enriches the feature representation of the model, thereby enhancing its generalization capabilities and category differentiation.
@@
Within the Contrastive Learning framework, we further investigated the impact of various data augmentation methods on model performance, as shown in Table VI, including Color Jitter, Resize, and Flip. The experimental results indicate that Resize and Flip have a minimal effect on model performance, showing almost no significant differences. In contrast, the inclusion of Color Jitter resulted in a noticeable improvement in model performance. This may be attributed to the fact that Resize and Flip do not substantially alter the overall distribution of the images, preventing Contrastive Learning from extracting valuable representation information. In contrast, Color Jitter modifies the color features of the images, disrupting their original distribution, which allows Contrastive Learning to better learn useful representation information from similar images. Even without employing further data augmentation techniques such as Resize and Flip, Color Jitter alone significantly enhances model performance.
TABLE VI: The effect of different augmentation methods, such as Collor Jitter, Gray Scale, Gaussian Blur, Resize and Flip on the performance of the model. We use the average IoU and average F1-score for evaluation, where all values are in % and higher values indicate better performance. SiamSeg w. Resize denotes the version of Siamseg with the Resize augment applied in contrastive learning. C.J. denotes Collor Jitter, G.S denotes Gray Scale, G.B. denotes Gaussian Blur. The task chosen in the table is POT IRRG → VAI IRRG (Potsdam IRRG to Vaihingen IRRG).
@@
SiamSeg w. All 65.55 77.07
IV-D4 Computational Complexity Analysis
In this study, we employed the MixTransformers architecture from Segformer [18] as our backbone network, which has a parameter count of 85.15M. This represents an approximately 30% increase compared to DeepLabV3+ [57], which has 62.7M parameters; however, this increase is accompanied by a significant improvement in performance. Although transformer-based backbone networks typically have a higher parameter count than traditional CNN-based networks, they offer the advantage of not requiring additional auxiliary networks, such as the complex architectures needed in adversarial learning methods.
The proposed method integrates Contrastive Learning with a shared-weight backbone, adding only two MLP heads (specifically, a projection head and a prediction head). Consequently, the increase in parameter count from this addition is minimal and does not impose a significant burden on overall computational complexity. In contrast, DualGAN [58] utilizes two generators and two discriminators, with each generator having 41.28M parameters and each discriminator 6.98M, resulting in a total parameter count of 97.56M—significantly higher than the proposed SiamSeg method.
Therefore, despite the slight increase in parameter count of the transformer-based SiamSeg, its efficient weight-sharing mechanism and lightweight MLP head design ensure that its overall computational complexity remains acceptable. Moreover, compared to complex architectures like DualGAN, it achieves a significant reduction in parameter count, demonstrating better computational efficiency.

### v3

@@ IV-D2 Visualization Results
In this section, we further visualize the model’s predictions to validate the outstanding performance of SiamSeg. Additionally, to explore the effectiveness of contrastive learning (C.L.), we visualize and compare the predictions of SiamSeg without C.L. (SiamSeg w./o. C.L.).
The overall segmentation results, as shown in Fig. 4, for AdaptSegNet, ProDA, and DualGAN, which are based on adversarial learning, are average due to the inability of these methods to be specially designed to account for the characteristics of a large domain gap in the RS domain. In contrast, the methods proposed by The CIA-UDA and DNT methods demonstrate excellent performance in RS image segmentation, indicating that the segmentation capabilities of a network can be significantly enhanced by addressing the challenge of inadequate learning in the target domain for RS cross-domain images. It is noteworthy that SiamSeg w./o. C.L. can achieve competitive results using only the ST method. However, due to the insufficient learning of the features in the target domain, there is considerable scope for improvement in performance. although SiamSeg w./o. C.L. achieves good segmentation results, focusing on the dotted box in the result graph, we can see that the prediction of the texture of the object and the correct classification are not as good as that of SiamSeg with the addition of contrastive learning, which shows that the addition of contrastive learning has reduced the domain gap between different domains. This also shows that the addition of contrastive learning reduces the domain gap between different domains by better learning the graph of the target domain.
@@
Figure 4: Visualization of results on Potsdam and Vaihingen datasets. The cross-domain tasks from top to bottom are Potsdam IRRG to Vaihingen IRRG and Potsdam RGB to Vaihingen IRRG. The categories represented by the different colors are listed at the bottom of the picture with their names and colors.
Figure 5: Visualization of results on LoveDA datasets. We conduct one task which is Rural to Urban. We provide the visualization results on LoveDA dataset. Since images in the testing dataset do not have annotations, we display the results of images in the validation dataset.
@@ IV-D3 Ablation Studies
Using the performance of SiamSeg and SiamSeg without Contrastive Learning (C.L.) in the task from Potsdam IRRG to Vaihingen IRRG, as shown in Table V, and visualization result in Fig. 5. We can clearly observe that SiamSeg outperforms SiamSeg w.o. C.L. across all categories after the introduction of C.L. This indicates that C.L. effectively enhances the model’s ability, in target domain, to perceive features across different categories. The additional supervisory signals provided by C.L. enable the model to learn image features more profoundly, significantly improving performance in cross-domain semantic segmentation tasks. The introduction of C.L. enriches the feature representation of the model, thereby enhancing its generalization capabilities and category differentiation. Solved the case of large domain gap in cross-domain RS image domains.
@@
Within the Contrastive Learning framework, we further investigated the impact of various data augmentation methods on model performance, as shown in Table VI, including Resize, Flip and Color Jitter. The experimental results indicate that Resize and Flip have a minimal effect on model performance, showing almost no significant differences. In contrast, the inclusion of Color Jitter resulted in a noticeable improvement in model performance. This may be attributed to the fact that Resize and Flip do not substantially alter the overall distribution of the images, preventing Contrastive Learning from extracting valuable representation information. In contrast, Color Jitter modifies the color features of the images, disrupting their original distribution, which allows Contrastive Learning to better learn useful representation information from similar images. Even without employing further data augmentation techniques such as Resize and Flip, Color Jitter alone significantly enhances model performance.
TABLE VI: The effect of different augmentation methods. SiamSeg w. Resize denotes the version of Siamseg with the Resize augment applied in contrastive learning. C.J. denotes Collor Jitter. The task chosen in the table is POT IRRG → VAI IRRG (Potsdam IRRG to Vaihingen IRRG).
@@
SiamSeg 65.55 77.07